OpenStudy (anonymous):

Miles continues on the climbing trail and ends up on point (-9, 6), The next bolt where he can clip his rope onto the rock is (-15, 8). Find the slope of the line that passes through these two points. (If the answer is a fraction, use the “/” on your keyboard. If the answer is undefined, write the word “undefined.”)

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 6-8 }{ -9+15 }=\frac{ -2 }{ 6}=?\]